OpenStudy (anonymous):

In 1980 median family income was about 17000 and in 2000 it was 43000 . What will the median family in be in 2006 if the slope of the like passing through 1980,17000 and 2000,43000 is 1300

OpenStudy (anonymous):

using points (0, 17000), (20, 43000), and (26, ?) with the slope given as 1300, and equation can be derived using the form y = mx + b. 17000 = 1300(0) + b gives you b = 17000. Thus the equation to use is y = 1300x + 17000. plug in 26 for the x component and you will have your median family income for year 2006.